The equation (x – 2)² + 1 = 2x – 3 is a
(a) linear equation
(b) quadratic equation
(c) cubic equation
(d) bi-quadratic equation

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Answer: b Explaination:Reason: We have (x – 2)² + 1 = 2x – 3 ⇒ x² + 4 – 2 × x × 2 + 1 = 2x – 3 ⇒ x² – 4x + 5 – 2x + 3 = 0 ∴ x² – 6x + 8 = 0, which is a quadratic equation.